Project A has cash flows of -$52,000, $26,500, $25,000, and $24,000 for years 0 to 3, respectively. Project B has an initial cost of -$50,000 and an annual cash inflow of $24,000 for three years. These are mutually exclusive projects. What is the crossover rate?
Answer #1

Incremental cash flow from year 1 =-52000+50000 =-2000
Cash flow year 1 =26500-24000=2500
Cash Flow year 2 =25000-24000 =1000
Cash Flow year 3 =24000-24000=0

Using financial calculator we calculate
CF0 =-2000; CF1 =2500; CF2 =1000; CF3 =0 CPT IRR =56.87%
Cross over rate =56.87%